# Transverse momentum asymmetry in the semi-inclusive electron positron annihilation process

### v1

@@ V Summary
Hadronization, being a non-perturbative QCD process, cannot be calculated from first principles and is typically described through phenomenological models. Within the framework of the QCD parton model, one can study the hadronization process through the behaviors of the single-hadron and the double-hadron or through FFs and diFFs. However, higher-twist FFs present significant challenges for determination due to power suppression. We notice that the transverse momentum of the produced jet or the intrinsic quark in semi-inclusive electron positron annihilation process exhibits twist-3 effects. Under this circumstance, we propose a novel observable: the transverse momentum asymmetry of the fragmenting quark with respect to the observed hadron direction. This asymmetry is defined as the difference between the transverse momentum components of the fragmenting quark relative to the observed hadron, specially in the $x$ and $y$ directions in the center-of-mass frame of leptons. It is a twist-3 effect and provides direct access to twist-3 TMD FFs.
@@
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grants No. 12405103, 12305106), and the Youth Innovation Technology Project of Higher School in Shandong Province (2023KJ146).
@@ Appendix A $\Xi^{(1)}_{\rho}$ contributions
Subsequently, we employ the relationships given in Eqs. (89)-(93) to express the functions $D_{d}$ and $G_{d}$ in terms of $D$ and $G$. This substitution leads us to the final result presented in Eq. (54).

### v2

@@ V Summary
Hadronization, being a nonperturbative QCD process, cannot be calculated from first principles and is typically described through phenomenological models. Within the framework of the QCD parton model, one can study the hadronization process through the behaviors of the single hadron and the double hadron or through FFs and diFFs. However, higher-twist FFs present significant challenges for determination due to power suppression. We notice that the transverse momentum of the produced jet or the intrinsic quark in the semi-inclusive electron positron annihilation process exhibits twist-3 effects. Under this circumstance, we propose a novel observable: the transverse momentum asymmetry of the fragmenting quark with respect to the observed hadron direction. This asymmetry is defined as the difference between the transverse momentum components of the fragmenting quark relative to the observed hadron, especially in the $x$ and $y$ directions in the center-of-mass frame of leptons. It is a twist-3 effect and provides direct access to twist-3 TMD FFs.
@@
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grants No. 12405103, and No. 12305106), and the Youth Innovation Technology Project of Higher School in Shandong Province (2023KJ146).
@@ Appendix A $\Xi^{(1)}_{\rho}$ CONTRIBUTIONS
Subsequently, we employ the relationships given in Eqs. (91)–(95) to express the functions $D_{d}$ and $G_{d}$ in terms of $D$ and $G$. This substitution leads us to the final result presented in Eq. (54).